5+(1/2-1/8)*12-8= n
Answer by tommyt3rd(5050)

n=5+(1/2-1/8)*12-8=
5+(4/8-1/8)*12-8=
5+(3/8)*12-8=

5+9/2-8=
10/2+9/2-8/2=
3/2
:)